Vertical Analysis help!Using the following data from the comparative balance sheet of Goody Company. December 31, 2017 December 31, 2016 Accounts receivable Inventory Total assets $418,320 816,720 2,490,000 $425,250 787,500 3,150,000 Illustrate vertical analysis. (Round percentages to 1 decimal place, e.g. 12.1%.) December 31, 2017 December 31, 2016 Goody Company Balance Sheets Amount Percentage Amount Percentage Accounts receivable $418,320 $816,720 $2,490,000 % $425,250 % $787,500 % $3,150,000 Inventory Total assets
Answer #1

Under vertical analysis;each item is represented as a % of total assets/Total liabilities and stockholders equity in the balance sheet.

December 31,2017 December 31,2016
AR 418320 (418320/2,490,000)=16.8% 425250 (425250/3,150,000)=13.5%
Inventory 816720 (816720/2,490,000)=32.8% 787500 (787500/3,150,000)=25%
Total assets 2,490,000 100% 3,150,000 100%